A square is inscribed in a circle whose equation is x^2+y^2=18. Find the dimension of the square.
Answer by stanbon(75887)

A square is inscribed in a circle whose equation is x^2+y^2=18. Find the dimension of the square.
The circle has radius = sqrt(18) = 3sqrt(2)
----
The square has diagonal = 2*3sqrt(2) = 6sqrt(2)
---------
Let each side of the equare be "x"::
x^2 + x^2 = (6sqrt(2))^2
-----
2x^2 = 72
x^2 = 36
x = side = 6
------
Dimensions of the square:: 6 by 6
-------------------
Cheers,
Stan H.
-------------